# Visual-auditory Extrinsic Contact Estimation

### v2

@@ Abstract
Keywords: Multimodal Perception, Contact-Rich Manipulation, Sim2Real
@@
footnotetext: * Denotes equal contribution.footnotetext: Correspondence to: {yixili, jayjun}@umich.edu.
1 Introduction
@@ 3 Methodology
The premise of our approach is that, while audio signatures vary across specific object interactions, they share consistent patterns that reflect the underlying contact geometry. A model trained on such data can learn to generalize by ignoring object-specific details and focusing on features indicative of contact type. For instance, point or line contacts tend to produce higher-pressure interactions, resulting in sharper, high-frequency sounds. In contrast, patch contacts generate a more diffuse sounds resembling white noise, with energy spread across a wider frequency range due to a larger contact area. A brief illustrative case study is provided in Appendix B using power spectral analysis.
@@
We further utilize the audio modality to acquire active feedback from extrinsic contact and interaction with the environment from the sine sweep impulse response that changes per contact mode type (refer to Appendix B), grasped object, and robot motion. The output of the model is a contact probability map that can be thresholded to obtain the contact patch as a single channel mask image.
4 Implementation
@@
Figure 4: Different contact modes learned from simulation data. The cropped depth is centered around the projected pixel coordinate of the EE pose. Sample predictions from test simulation data are shown as contact probability maps overlaid on top of scene depth.
Training Details: We train our model using the pixel-wise binary cross entropy loss with logits using the ground-truth contact masks obtained from simulation. We train for 20 epochs with a learning rate of $45e-$ 4 on a single A6000 GPU.
@@
Additionally, for collecting the real-world depth images, we primarily use the monocular metric depth estimation model Depth-Pro [29] to generate hole-less depth map from raw RGB images of the scene (refer to Appendix A to see noisy depth images that are significantly out-of-distribution). We collected data in 4 cases: general, occlusion, near-contact, and different surface cases. We use RAFT [30] to generate optical flow images. We use the same audio collection pipleline as audio collection for test data collection.
@@
5 Experiments
Models Binary Contact Detection Contact Location, Shape, & Size
Prsn $\uparrow$ Rcll $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ BCE $\downarrow$ IOU $\uparrow$ CD ($\text{px}^{2}$) $\downarrow$
Im2Contact 0.70 0.94 0.81 0.69 0.023 0.058 19.2
VA2Contact w/o flow 0.98 0.95 0.96 0.95 0.030 0.073 13.6
VA2Contact 0.86 0.90 0.88 0.83 0.027 0.28 4.8
Table 1: Real world results for sim-to-real extrinsic contact estimation on S1 to test on general contact cases. Models are evaluated on binary contact detection and the accuracy of contact geometry estimation, which is only provided for the true positive cases where the model predicts a mask above a threshold level and the ground-truth is in-contact. CD: Chamfer Distance. IOU: Intersection over Union. BCE: Binary cross-entropy.
5.1 Extrinsic Contact Estimation
Baselines and Ablations: We compare our visual-auditory method against a similar model-free and vision-only method Im2Contact [1]. We also evaluate a variant of VA2Contact trained without optical flow to experiment if audio can fully replace it.
@@ 5.1 Extrinsic Contact Estimation
  • S1: Generalization to unseen gripper-held object types and geometry. (Table 1 and Figure 9)
@@
  • S2: Generalization to unseen table surface to introduce audio variation. (Appendix C)
  • S3: Near contact cases that are not in contact. (Appendix C)
@@
  • S4: Cases where the contact between gripper-held object and the table is occluded. (Appendix C)
@@
Experimental Results: The main results of the sim2real transfer of visual-auditory extrinsic contact estimation in the real world setup are shown in Table 1 and Figure 9. We further explore applications to edge cases such as visual occlusions of the contact and near-contact setups where the active-acoustic signal could provide important cues. In the general case, the model with audio overall exhibits higher recall and F1 score, indicating a superior ability to detect true contacts. This suggests that audio cues help reduce false negatives without substantially increasing false positives. VA2Contact w/o optical flow scores the highest across all metrics for binary contact detection because audio provides a more direct information about contact whereas optical flow just tells us where things moved, which might not be directly correlated. However, in terms of the actual contact patch prediction, VA2Contact which uses both active-audio and optical flow performs the best as shown in Figure 9. For other scenarios (S2-4) where an additional sensor modality can be useful, we observe that VA2Contact does improve contact perception under heavy visual ambiguities. For more details and predictions, refer to Appendix C.
5.2 Contact-aware Policy Learning
Figure 6: Real-world wiping task scene with contact-aware diffusion policy, enabled by VA2Contact.
@@ 5.2 Contact-aware Policy Learning
Given this sim2real visual-auditory extrinsic contact sensing capability, we evaluate whether this can improve the downstream policy learning and performance for contact-rich and occlusion-heavy manipulation tasks. We selected a wiping task, which requires the robot to hold a marker eraser and maintain contact to erase the markers. To test this, we train a baseline diffusion policy with just an RGB image stream and a contact-aware diffusion policy as in Figure 2 that takes the contact probability map image in addition to the RGB image stream as inputs. We evaluate the performances over 10 rollouts for each model. The baseline model achieved 4/10 success due to inconsistent contact with the whiteboard whereas the contact-aware policy achieved 8/10 success rate.
@@
6 Conclusion and Discussions
Summary: We presented a method trained entirely in simulation that accurately estimates extrinsic contacts by leveraging multimodal signals—fingertip audio, vision, and proprioception—without relying on prior knowledge of the object or environment. Active audio sensing is shown to provide rich, local information about contact events and modes, effectively complementing vision in contact-rich manipulation tasks. By injecting realistic audio signals into simulation data, we bypass the difficulty of simulating audio while still improving contact prediction performance. Our method demonstrates strong sim-to-real transfer, especially in visually ambiguous contact scenarios, outperforming baselines that exclude audio input.
Limitations: Despite its effectiveness, our system faces several limitations. Internal vibrations from the robot—especially within the Franka gripper—introduce significant noise, reducing the signal-to-noise ratio (SNR) and impacting audio quality. Improving vibration isolation at the fingertip could substantially enhance signal clarity and sensing robustness.
The contact masks in the real world test dataset are annotated manually, with the help of RGB image collected along with depth. It is noted that these masks are subject to the annotator’s subjective experience due to the occlusion and difficulty to directly observe the ground truth contact.
The current audio representation, while effective, may not capture the full range of contact features. Exploring alternative encodings or integrating large pretrained audio models could yield more expressive and robust features for contact reasoning.
Our dataset focuses on rigid objects with low vibration absorption to ensure clean audio, but this restricts generalizability. Softer, more deformable objects and liquid pose challenges due to annotation complexity and high signal variability, yet incorporating them is essential for broader applicability.
In addition, we have yet to fully explore contact chains, where signals travel across multiple interfaces and suffer cumulative loss. Addressing these scenarios will require more sensitive sensing, improved actuation—such as the potential use of vibration motors—and careful attention to SNR across all system components.
Appendix A Real World Depth Images
The figure below displays how noisy depth images can be despite using a higher-end high resolution 3D cameras compared to a simulation environment that returns an optimal depth maps.
Figure 7: A comparison between a depth image from simulation and a raw and a sample processed depth images from Photoneo 3D scanner. The real depth images present a significant distribution shift from the simulation training data.
Appendix B Power Spectral Analysis of Acoustic Signatures across Contact Modes
As illustrated in Figure 8, each contact mode exhibits distinct characteristics in its power spectral density (PSD), revealing informative patterns tied to contact geometry. In the low-frequency range (0–100 Hz), all modes show a steep initial drop in spectral power, with the patch contact displaying the lowest overall magnitude—indicative of more dampened or lower-energy vibrations. The first peak, observed around 100–300 Hz, is both wider and higher for the point contact compared to line and patch contacts. This reflects the sharper, higher-energy acoustic responses produced by the small contact area and high contact pressure of point interactions. In the mid-frequency range (300–1000 Hz), the line contact maintains higher PSD values than the point contact, which itself remains consistently above the patch. This suggests that line contacts produce more sustained vibrational energy across a broader range of frequencies. In the high-frequency region beyond 1000 Hz, all contact types show a gradual convergence, though the line contact still dominates in PSD magnitude, and the patch contact continues to exhibit the lowest profile. Overall, these trends indicate that the size and type of contact area—point, line, or patch—have a significant impact on the spectral shape of the resulting acoustic signal. The presence of consistent and structured spectral differences highlights the potential for learning-based models to classify contact modes based on audio features alone.
As our method utilizes the contact mode type labels to inject audio into the simulated episodes of contact, we analyze the frequency components of the different contact modes using various household objects and different surface types to showcase how each mode displays similar features despite variances in vibrational properties, material type, rigidity, and content. We recorded 3s long audio waveforms where a sliding contact is made between each object and surface of the environment. However, it is noted that it is difficult to exert a consistent force over different setups so we collect three audio data per setup and average. We collected the audio signals over four object types (a box, an empty plastic bottle, a power strip, a metallic screw driver set case) on two different surface types (a powder coated steel and a wooden surface) and analyzed the power spectral density averaged across different contact modes. The acoustic signal features and patterns can be explained through the contact pressure and the contact size it makes with the environment. Assuming the same force is being exerted, the contact pressure is highest from the point, line, and patch, in this order as the contact area increases. The patch contact in general has a relatively horizontal profile after the steep drop in the low frequency range from 0 to 100 Hz. In contrast, shifting focus to the first peak of the point and line contact plots, we see slight difference in the bandwidth of the peak where the point contact is wider and higher. As the frequency increases the point contact plot maintains a profile that stays under the line contact’s profile. It should be noted that the forced applied at different contact modes is not consistent and there are more variety in combinations possible in the real world scenarios. However, this figure presents that there exists a learnable pattern in which audio signal may be from which contact mode.
Figure 8: Averaged power spectral density (PSD) of acoustic feedback from sliding contact interactions across three contact modes: point, line, and patch. Point contact exhibits a broad, high-amplitude peak in the 100–300 Hz range, line contact shows elevated energy across mid to high frequencies, and patch contact presents a flatter, more damped spectral profile. Results are averaged over four objects and two surface types, highlighting the consistent, mode-specific spectral signatures in contact-generated audio.
Appendix C More Results for Extrinsic Contact Estimation
Different Surface Type Cases (S2): For scenarios involving a different surface in Table 2, both models scores high in binary contact detection metrics. For the contact geometry estimation, the performance of the two models are similar for the model with audio or even better for the model without audio in terms of IOU. This indicates that while our method can transfer to different surface types, which might lead to a variation in audio feedback. The model without audio tends to be less affected by this change as it solely reasons through vision. This indicates that while visual cues are sufficient for high precision, audio input can be generalized across varying surfaces with consistent contact detection.
Models Binary Contact Detection Contact Location, Shape, & Size
Prsn $\uparrow$ Rcll $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ BCE $\downarrow$ IOU $\uparrow$ CD ($\text{px}^{2}$) $\downarrow$
Im2Contact 1.00 0.78 0.88 0.78 0.04 0.139 10.52
VA2Contact w/o flow 1.00 0.86 0.92 0.86 0.060 0.116 10.74
VA2Contact 1.00 0.94 0.97 0.94 0.054 0.221 4.36
Table 2: Real world results for sim-to-real extrinsic contact estimation on S2 for Evaluating on Different Surface Types. Models are evaluated on binary contact detection and the accuracy of contact geometry estimation, which is only provided for the true positive cases where the model predicts a mask above a threshold level and the ground-truth is in-contact. CD: Chamfer Distance. IOU: Intersection over Union. BCE: Binary cross-entropy.
Near-contact Cases (S3): In the near contact scenario in Table 3, where no ground truth contact masks are available as there is no contact event, models that have audio significantly outperform the model that relies solely on vision. Most of Im2Contact’s predictions are false positives, which indicates the inability to disambiguate near contact cases. Yet our VA2Contact can effectively probe with active-audio sensing to differentiate being in contact and not in contact despite visual ambiguity.
Models Binary Contact Detection
TP $\uparrow$ TN $\uparrow$ FP $\downarrow$ FN $\downarrow$
Im2Contact 0 2 48 0
VA2Contact w/o flow 0 48 2 0
VA2Contact 0 49 1 0
Table 3: Real world results for sim-to-real extrinsic contact estimation on S3 for Evaluating on Near-Contact Cases where there is no actual contact being made. Models are evaluated on binary contact detection with TP (true positive), TN (true negative), FP (false positive), FN (false negative).
Occlusion Cases (S4): Under occlusion in Table 4, the benefits of audio are even more pronounced that the VA2Contact with audio outperforms Im2Contact. Our model with audio achieves a recall of 0.94 compared to 0.78 without audio, which indicates VA2Contact’s ability to better perceive contacts under occlusions, and an F1 score of 0.97 versus 0.88. The accuracy also improves significantly (0.94 vs. 0.78). These improvements highlight the model’s enhanced capability to detect contacts that are visually obstructed, leveraging audio cues to compensate for the lack of visual information.
Models Binary Contact Detection Contact Location, Shape, & Size
Prsn $\uparrow$ Rcll $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ BCE $\downarrow$ IOU $\uparrow$ CD ($\text{px}^{2}$) $\downarrow$
Im2Contact 0.91 0.89 0.90 0.82 0.029 0.079 15.3
VA2Contact w/o flow 1.00 0.87 0.93 0.88 0.025 0.099 10.84
VA2Contact 0.96 1.00 0.98 0.96 0.034 0.22 10.16
Table 4: Real world results for sim-to-real extrinsic contact estimation on S4 for Occlusions. Models are evaluated on binary contact detection and the accuracy of contact geometry estimation, which is only provided for the true positive cases where the model predicts a mask above a threshold level and the ground-truth is in-contact. CD: Chamfer Distance. IOU: Intersection over Union. BCE: Binary cross-entropy.

### v3

@@ Abstract
I Introduction
@@ III Methodology
The premise of our approach is that, while audio signatures vary across specific object interactions, they share consistent patterns that reflect the underlying contact geometry. A model trained on such data can learn to generalize by ignoring object-specific details and focusing on features indicative of contact type. For instance, point or line contacts tend to produce higher-pressure interactions, resulting in sharper, high-frequency sounds. In contrast, patch contacts generate a more diffuse sounds resembling white noise, with energy spread across a wider frequency range due to a larger contact area.
@@
We further utilize the audio modality to acquire active feedback from extrinsic contact and interaction with the environment from the sine sweep impulse response that changes per contact mode type, grasped object, and robot motion. The output of the model is a contact probability map that can be thresholded to obtain the contact patch as a single channel mask image.
IV Implementation
@@
Figure 4: Different contact modes learned from simulation data. The cropped depth is centered around the projected pixel coordinate of the $E$ pose. Sample predictions from test simulation data are shown as contact probability maps overlaid on top of scene depth.
Training Details: We train our model with pixel-wise binary cross entropy loss with logits using the ground-truth contact masks obtained from simulation. We train for 20 epochs with a learning rate of 5e45e-4 on a single A6000 GPU.
@@
Additionally, for collecting the real-world depth images, we primarily use the monocular metric depth estimation model Depth-Pro [29] to generate hole-less depth map from raw RGB images of the scene. We collected data in 4 cases: general, occlusion, near-contact, and different surface cases. We use RAFT [30] to generate optical flow images. We use the same audio collection pipeline as audio collection for test data collection.
@@
V Experiments
Models Binary Contact Detection (macro) Contact Location, Shape, & Size (macro)
Prsn $\uparrow$ Rcll $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ BCE $\downarrow$ IOU $\uparrow$ CD ($\mathrm{px}^{2}$) $\downarrow$
Im2Contact 0.87 0.87 0.86 0.58 0.031 0.092 15.01
VA2Contact w/o flow 0.99 0.89 0.94 0.91 0.038 0.096 11.73
VA2Contact 0.94 0.95 0.94 0.93 0.038 0.240 6.44
TABLE I: Real world results for sim-to-real extrinsic contact estimation on all 4 scenarios. Models are evaluated on binary contact detection and the accuracy of contact geometry estimation, which is only provided for the true positive cases where the model predicts a mask above a threshold level and the ground-truth is in-contact. CD: Chamfer Distance. IOU: Intersection over Union. BCE: Binary cross-entropy.
Subset Model Prsn $\uparrow$ Rcll $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ BCE $\downarrow$ IOU $\uparrow$ CD ($\mathrm{px}^{2}$) $\downarrow$
S2: Surface Types Im2Contact 1.00 0.78 0.88 0.78 0.040 0.139 10.52
VA2Contact w/o flow 1.00 0.86 0.92 0.86 0.060 0.116 10.74
VA2Contact 1.00 0.94 0.97 0.94 0.054 0.221 4.36
S3: Near Contact (no positive GT; Precision/Recall/F1 not defined)
Im2Contact TN=2,  FP=48 0.04
VA2Contact w/o flow TN=48,  FP=2 0.96
VA2Contact TN=49,  FP=1 0.98
S4: Occlusions Im2Contact 0.91 0.89 0.90 0.82 0.029 0.079 15.30
VA2Contact w/o flow 1.00 0.87 0.93 0.88 0.025 0.099 10.84
VA2Contact 0.96 1.00 0.98 0.96 0.034 0.220 10.16
TABLE II: Real-world results for sim-to-real extrinsic contact estimation. Binary detection metrics (Precision/Recall/F1/Accuracy) are reported where positive ground-truth contacts exist (S2, S4). For S3, only near-contact negatives are present; we therefore report true negatives (TN), false positives (FP), and accuracy. Geometry metrics (BCE, IOU, Chamfer Distance) are evaluated on true positives only.00footnotetext: Precision, Recall, and F1 are undefined when no positive ground-truth contacts occur.
V-A Extrinsic Contact Estimation
Baselines and Ablations: We compare our visual-auditory method against a similar model-free and vision-only method Im2Contact [1]. We also evaluate a variant of VA2Contact trained without optical flow to experiment if audio can fully replace it. Simulation Test Scenarios: We first evaluate our methods in simulation
@@ V-A Extrinsic Contact Estimation
  • S1: Generalization to unseen gripper-held object types and geometry. (Figure 5)
@@
  • S2: Generalization to unseen table surface to introduce audio variation. (Figure 6)
  • S3: Near contact cases that are not in contact.
@@
  • S4: Cases where the contact between gripper-held object and the table is occluded. (Figure 6)
In total, we collected 310 evaluation samples. The number per contact mode is free:point:line:patch = {48:37:41:34} for S1, {0:15:15:20} for S2, {4:14:9:23} for S3, and {50:0:0:0} for S4.
@@
Experimental Results: The main results of the sim2real transfer of visual-auditory extrinsic contact estimation in the real world setup are shown in Table I and Figure 5. Here we show the performance of our methods across all 4 scenarios. In the general case, the model with audio overall exhibits higher recall and F1 score, indicating a superior ability to detect true contacts. This suggests that audio cues help reduce false negatives without substantially increasing false positives. VA2Contact w/o optical flow scores the highest across all metrics for binary contact detection because audio provides a more direct information about contact whereas optical flow just tells us where things moved, which might not be directly correlated. However, in terms of the actual contact patch prediction, VA2Contact which uses both active-audio and optical flow performs the best as shown in Figure 5.
We also explore edge cases such as visual occlusions and near-contact scenarios where active-acoustic signals provide critical cues, as shown in Figure 6. We evaluate VA2Contact under three challenging real-world settings: different surface types, near-contact, and occlusions. On different surfaces (S2, Table II), both vision-only and audio-augmented models achieve strong binary contact detection. While IOU is slightly better without audio, our method generalizes well despite variation in acoustic feedback, showing audio input remains effective under domain shifts. In near-contact scenarios (S3, Table II), VA2Contact significantly reduces false positives compared to Im2Contact, indicating its strength in disambiguating ambiguous visual cases via active sensing. Under occlusions (S4, Table II), VA2Contact consistently outperforms all baselines in both detection and geometry metrics (e.g., Recall: 0.94 vs. 0.78; F1: 0.97 vs. 0.88), demonstrating that audio cues robustly compensate for missing visual information.
V-B Contact-aware Policy Learning
@@
Given this sim2real visual-auditory extrinsic contact sensing capability, we evaluate whether this can improve the downstream policy learning and performance for contact-rich and occlusion-heavy manipulation tasks. We selected a wiping task, which requires the robot to hold a marker eraser and maintain contact to erase the markers, as shown in Figure 7. To test this, we train a baseline diffusion policy with just an RGB image stream and a contact-aware diffusion policy as in Figure 2 that takes the contact probability map image in addition to the RGB image stream as inputs. We evaluate the performances over 10 rollouts for each model. The baseline model achieved 4/10 success due to inconsistent contact with the whiteboard whereas the contact-aware policy achieved 8/10 success rate.
Figure 7: Real-world wiping task using a contact-aware diffusion policy enabled by VA2Contact. (a) Experimental setup where the robot wipes a whiteboard using our method. (b, c) Comparison of baseline and VA2Contact wiping a drawn face. (d, e) Comparison of baseline and VA2Contact wiping a boat. In both tasks, VA2Contact enables more consistent and complete contact with the surface, leading to better task performance.
@@
VI Conclusion and Discussions
This paper presented VA2Contact, a simulation-trained method for estimating extrinsic contacts using multimodal inputs: fingertip audio, vision, and proprioception. Active audio sensing provides local cues that complement vision, enabling robust contact perception under occlusions and ambiguity. To bypass the challenge of simulating audio, we introduce a real-to-sim hallucination strategy that injects real audio into training, enabling strong zero-shot transfer to the real world and outperforming visual-only baselines.
Despite promising results, limitations remain. Internal robot vibrations, especially from the Franka gripper, reduce audio signal quality. Manual annotations of contact masks may be biased due to occlusion. Our current audio representation may not fully capture rich contact cues—future work may explore pretrained models. Finally, the dataset focuses on rigid objects; generalizing to deformable or liquid-filled items and contact chains will require better sensing and annotation strategies.